OpenStudy (anonymous):

find the function f(x)= x+(1/x), g(x)=x+14/x+2 (fog)(x)= (gof)(x)= (fof)(x)= (gog)(x)=

OpenStudy (anonymous):

what does "fog" mean?

OpenStudy (abb0t):

f(x) composed of g(x).

OpenStudy (anonymous):

right that means f(g(x))

OpenStudy (anonymous):

That means that for all the "x" f(x) we replace with g(x) but we know that g(x)= x+14/x+2 so we replace with that instead.

OpenStudy (abb0t):

ohhhhh. i get it now!!

OpenStudy (anonymous):

\[so \frac{ x+14 }{ x+2 } + \frac{ 1 }{ x+14/x+2} ?\]

OpenStudy (anonymous):

yeah that's basically it. You can simplify it and rewrite it if your teacher told you to.

OpenStudy (anonymous):

(gof)(x) is the other way around. That means that for all the "x" g(x) we replace with f(x)

OpenStudy (anonymous):

could you give me a hand with simplifying the g(f(x))

OpenStudy (anonymous):

I don't think we can simplify it even further. :/